.


A Jollibee delivery man travels 5 km, North and then 3 km, East to deliver the order. What is angle from the + x- axis is the delivery man?

1
Expert's answer
2022-03-17T09:35:37-0400
"\theta=\tan^{-1}\frac{d_y}{d_x}=\tan^{-1}\frac{5}{3}=59^\circ"